OpenStudy (anonymous):

Factor by grouping. 3x^2 + 7x – 6 A. (3x – 2)(x – 3) B. (3x – 2)(x + 3) C. (x + 3)(3x + 2) D. (3x + 2)(x – 3)

OpenStudy (anonymous):

is anybody there?

OpenStudy (anonymous):

r u sure this r the options? no of them seem to fit

OpenStudy (anonymous):

yes I go it. thanks though

OpenStudy (anonymous):

oh ook

OpenStudy (anonymous):

can you check my next answer?

OpenStudy (anonymous):

Solve the system using elimination. 3x + 3y = –9 3x – 3y = 21 A. (3, –6) B. (–5, 2) C. (3, 3) D. (2, –5)

OpenStudy (anonymous):

@doubledrive there is an answer for the first one...

OpenStudy (anonymous):

is D (2,-5) calcmathlete

OpenStudy (anonymous):

3x + 3y = -9 3x - 3y = 21 ______________ Just add them, 6x + 0y = 12 6x = 12 x = 2 Now that you have x, which is 2, plug 2 back in one of the equations. 3(2) + 3y = -9 6 + 3y = -9 -6 -6 Subtract six from both sides. 3y = -15 y = -5 Now you have both x and y, just write them both as coordinates, (2,-5)

OpenStudy (anonymous):

I was referring to the first question on here. THe original.

OpenStudy (anonymous):

@Calcmathlete, chadandtricia_2009 already figure out the first question

OpenStudy (anonymous):

Oh ok. I didn't see that. I was saying that because I saw that you there wan't an answer for it.

OpenStudy (anonymous):

can you help me on the question I have posted Calcmathlete?